A circular loop is placed in uniform magnetic field of 1 Tesla as shown. If the radius of loop is 1m the magnetic force on loop is

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
There is no image attached. Please check and repost the question with an attachment. I will be happy to help you.
Khimraj
4 years ago
ϕ=πr2B dϕdt=2πr.dbdt when r=2×10−2manddrdt=10−3m/s e=dϕdt=2×227×(2×10−2)(0.02)×10−3 =2.5×10−6V =2.5μV